# Meta-Learning Guided Pruning for Few-Shot Plant Pathology on Edge Devices

### v1

@@ IV-R Multi-Resolution Robustness Analysis
TABLE XIX: Resolution Robustness (5-Way 5-Shot)
@@
IV-S Deployment Efficiency Metrics
@@
TABLE $X$: Deployment Efficiency Score (DES)
@@
TABLE XXI: Cross-Platform Inference Efficiency
@@ V-D Random Seed Analysis
TABLE XXII: Random Seed Impact Analysis (5-Way 5-Shot, 30% params)
@@ VI-A2 Component Contributions
  • Layer-Adaptive Thresholds: Contributes 2.8% improvement by respecting the hierarchical role of different network depths.
@@ VI-B Deployment Readiness
    Deployment Efficiency Score: DES metric reveals 4.7× improvement over ProtoNet baseline, simultaneously optimizing accuracy, FPS, model size, and energy—a holistic measure of deployment readiness.
@@ VII Conclusion: Validated Framework for Practical Deployment
This work presents a comprehensively validated framework for deploying few-shot plant disease detection on resource-constrained edge devices. The Disease-Aware Channel Importance Scoring (DACIS) mechanism and three-stage Prune-then-Meta-Learn-then-Prune (PMP) pipeline synergistically combine neural network compression with few-shot meta-learning, achieving substantial improvements in both accuracy and deployment efficiency.
@@ 3. Empirical Validation: Comprehensive experiments across multiple protocols yield strong evidence:
4. Deployment-Ready Systems: Introduction of deployment-focused metrics (DES, FSI, CSG) and evaluation protocols (simulated temporal generalization, multi-resolution, severity stratification) that better capture real-world deployment constraints than standard benchmarks.
@@ VII-C Future Research Directions
  • Hardware-Aware NAS: Co-optimization of network architecture and pruning strategy for specific embedded hardware (ARM processors, TPUs, quantum accelerators).
@@ VII-E Framework Extensibility
  • Meta-Learning Backend: Any gradient-based meta-learning algorithm (MAML, Reptile, Meta-SGD) can replace ProtoNet as the base learner.

### v2

@@ IV-L2 Multi-Resolution and Resolution Robustness Analysis
TABLE XIX: Resolution Robustness (5-Way 5-Shot). Inference at resolutions 64-448 pixels (trained at 224×224). Drop = accuracy degradation from peak. DACIS demonstrates superior stability across resolutions versus magnitude pruning and baseline.
@@
IV-M Deployment Efficiency Metrics
@@
TABLE $X$: Deployment Efficiency Score (DES) Analysis. DES = (Accuracy × Compression) / Latency, combining classification performance, compression ratio, and Raspberry Pi inference time. Scores for baseline, magnitude-pruned, and DACIS-pruned models.
@@
TABLE XXI: Cross-Platform Inference Efficiency (30% compression). Latency measured on GPU (NVIDIA RTX 3090), CPU (Intel i7), and Edge (Raspberry Pi 4). DES = (Accuracy $\times$ Compression) / Latency metric.
@@ V-D Random Seed Analysis
TABLE XXII: Random Seed Impact Analysis (5-Way 5-Shot, 30% params). Evaluates PMP-DACIS robustness across 10 random initializations on PlantVillage. Low variance confirms reproducible and stable performance.
@@ VI-A2 Component Contributions
  • Layer-Adaptive Thresholds: Contribute 2.8% improvement by respecting the hierarchical role of different network depths.
@@ VI-B Deployment Readiness
    Deployment Efficiency: DES reveals a 4.7× improvement over the ProtoNet baseline, simultaneously optimizing accuracy, FPS, model size, and energy—a holistic measure of deployment readiness.
@@ VII Conclusion: Validated Framework for Practical Deployment
This work presents a comprehensively validated framework for deploying few-shot plant disease detection on resource-constrained edge devices. The DACIS mechanism and three-stage PMP pipeline synergistically combine neural network compression with few-shot meta-learning, achieving substantial improvements in both accuracy and deployment efficiency.
@@ 3. Empirical Validation: Comprehensive experiments across multiple protocols yield strong evidence:
4. Deployment-Ready Systems: Introduction of deployment-focused metrics (Deployment Efficiency Score, Feature Stability Index, Cross-Stage Generalization) and evaluation protocols (simulated temporal generalization, multi-resolution, severity stratification) that better capture real-world deployment constraints than standard benchmarks.
@@ VII-C Future Research Directions
  • Hardware-Aware Neural Architecture Search: Co-optimization of network architecture and pruning strategy for specific embedded hardware (ARM processors, TPUs, quantum accelerators).
@@ VII-E Framework Extensibility
  • Meta-Learning Backend: Any gradient-based meta-learning algorithm (MAML, Reptile, Meta-Stochastic Gradient Descent) can replace ProtoNet as the base learner.
@@
Acknowledgments and Funding
This research was supported by the Department of Computer Science and Engineering at Muffakham Jah College of Engineering and Technology. The authors acknowledge computational resources provided by the institution’s high-performance computing facilities for model training and evaluation.